OpenStudy (anonymous):

how would you put 2x=10+5y in mx+b form?

OpenStudy (anonymous):

2x=10+5y 2x-5y=10 -5y=-2x+10 y=(-2x+10)/-5 y=2x/5-2

OpenStudy (anonymous):

how do u do this?

OpenStudy (anonymous):

You just need to rearrange the equation so y is by itself :)